# Distributed Evolution Strategies with Multi-Level Learning for Large-Scale Black-Box Optimization

### v1

@@ Abstract
In the post-Moore era, the main performance gains of black-box optimizers are increasingly depending upon parallelism, especially for large-scale optimization (LSO). In this paper, we propose to parallelize the well-established covariance matrix adaptation evolution strategy (CMA-ES) and in particular its one latest variant called limited-memory CMA (LM-CMA) for LSO. To achieve scalability while maintaining the invariance property as much as possible, we present a multilevel learning-based meta-framework. Owing to its hierarchically organized structure, Meta-ES is well-suited to implement our distributed meta-framework, wherein the outer-ES controls strategy parameters while all parallel inner-ESs run the serial LM-CMA with different settings. For the distribution mean update of the outer-ES, both the elitist and multi-recombination strategy are used in parallel to avoid stagnation and regression, respectively. To exploit spatiotemporal information, the global step-size adaptation combines Meta-ES with the parallel cumulative step-size adaptation. After each isolation time, our meta-framework employs both the structure and parameter learning strategy to combine aligned evolution paths for CMA reconstruction. Experiments on a set of large-scale benchmarking functions with memory-intensive evaluations, arguably reflecting many data-driven optimization problems, validate the benefits (e.g., scalability w.r.t. CPU cores, effectiveness w.r.t. solution quality, and adaptability w.r.t. second-order learning) and costs of our meta-framework.
@@
Specifically, we consider the derandomized evolution strategy with covariance matrix adaptation (CMA-ES [16, 17, 18]) and in particular its one latest variant called limited-memory CMA (LM-CMA [19, 20]) for large-scale black-box optimization (BBO). As stated in the popular Nature review [12], “CMA-ES is widely regarded as (one of) the state of the art in numerical (black-box) optimization” with competitive performance on many benchmarking functions [21, 22, 23, 24, 25] and challenging applications (such as [26, 27, 28, 29, 30, 31], just to name a few). Although typically its absolute runtime is of minor relevance [17] for low-dimensional (e.g., 50) cases, it cannot be ignored in the distributed computing context for large-scale (e.g., 1000) optimization (LSO) because of its (at least) quadratic complexity w.r.t. each sampling. Since its limited-memory variant (LM-CMA) can fit more for memory hierarchy and distributed communication, our aim is to extend LM-CMA to the modern cloud/clustering computing environment for LSO, in order to show effectiveness and scalability (an essential feature for any distributed algorithms [32]) while approximating the attractive invariance property of CMA-ES [33] as much as possible.
@@
In their seminal paper [17], Hansen and Ostermeier pointed out four fundamental demands for ESs: adaptation, performance, invariance [34], and stationarity (i.e., unbiasedness under random selection). Clearly, these demands111As previously stated in the classic ES review by Beyer and Schwefel [35], if some principle is deviated when we design the ES-based variant, the designed optimizer often needs to be well-tested. are also highly desirable for any distributed ES (DES), in order to obtain efficiency and generalization/transferability [36]. To meet these demands in the distributed computing environment, we adopt the multilevel learning perspective for evolution (recently published in PNAS [37, 38]) to model and design the scalable DES framework, wherein Meta-ES [39, 40] and LM-CMA could be naturally combined together to enjoy the best of both worlds. Three main contributions of our paper to DES for large-scale BBO are presented as below:
1) We first analyze the parallelism opportunities and challenges of two different ES families (i.e., CMA-ES and Meta-ES) under two common models (i.e., master-salve and island). See Section II for details. For CMA-ES, efficient (distributed shared) memory usage is a prerequisite to design its distributed algorithm, and instead its large-scale variants are preferred as the baseline computing unit because of their much lower (e.g., ) complexity. For Meta-ES, the choice of its inner-ES loop significantly impacts its convergence rate, since the fixed global step-size adaptation used in its inner-ESs may be not optimal for many challenging problems when a long isolation time is used for population diversity as well as low communication costs. We argue that multilevel learning for biological evolution (MLE) is a natural way to hierarchically combine LM-CMA with Meta-ES under distributed computing (see Section III for details).
@@
2) Inspired by the recent evolution theory (i.e., MLE) [37], we propose a multilevel learning-based meta-framework for DES to exploit spatio-temporal information (if available) on-the-fly to accelerate global convergence while maintaining meta-population diversity under the affordable cloud/clustering computing platform. Within this meta-framework, four following key design choices for DES are made reasonably, in order to balance search efficiency (w.r.t. convergence rate, absolute runtime, and diversity) and extra computing cost brought by distributed enhancements (e.g., distributed scheduling, load balancing, fault-tolerance on computing nodes, and data exchanges over the network, etc.) [41].
First, owing to its hierarchically organized structure, Meta-ES is well-suited to implement the multilevel meta-framework for DES. Generally, the (unique) outer-ES controls only strategy parameters, which parameterize the (typically Gaussian) search/mutation distribution [42], while parallel inner-ESs evolve independent sub-populations (with different strategy parameters initialized via the outer-ES level) within each isolation time. Unlike all of previous Meta-ESs which employ fixed strategy parameters for inner-ESs, here we use the computationally efficient LM-CMA as the base to learn the (local) topology, some of which are then collected to approximate the global topology at the outer-ES level. Refer to Section III-A for details.
Second, at the outer-ES level, the mean update of search distribution may suffer from stagnation when the elitist strategy from previous Meta-ESs is used (to our knowledge, this finding is dated back to Schwefel [43]). If only the multi-recombination strategy is used [44], it may suffer from the regression issue on some functions (e.g., Cigar). In our meta-framework, both the elitist and multi-recombination strategy are used in a parallel fashion, to enjoy the better of two worlds (see Section III-B for details).
Third, the global step-size (also called mutation strength) self-adaptation for DES combines the less-known Meta-ES strategy with the well-known cumulative step-size adaptation (CSA), in order to exploit both the spatial and temporal (non-local) information. As stated previously by Rudolph [45], the optimization/control problem of strategy parameters is multi-modal [17] and noisy (from sampling). Although they sometimes can be seen as hyper-parameters when an offline perspective is used [46, 47], strategy parameters themselves are usually controlled in an online fashion during evolution, in order to timely utilize (possibly) topology information to maximize convergence progress (see Section III-C for details).
Fourth, to keep a sensible trade-off among efficiency, stability, and scalability for DES, we extend the collective learning strategy (originally proposed by Schwefel [48]) to the distributed computing scenario mainly from the following two aspects: i) after each isolation time all evolution paths from parallel LM-CMAs will be aggregated (via weighted averaging at the outer-ES level) as the shared baseline to reconstruct initial covariance matrix for inner-ESs (i.e., LM-CMAs) of the next isolation time; ii) different LM-CMA structures, simply parameterized as the number of evolution paths for CMA reconstruction, are adapted at the outer-ES level, in order to approximate the global geometry of the fitness landscape as soon as possible (see Section III-D for details).
3) To validate the effectiveness and scalability [49] of our proposed meta-framework, we conduct numerical experiments on a large set of large-scale (i.e., 2000-d) BBO functions. The standard rotation operation used to avoid separability results in memory-expensive fitness evaluations, arguably reflecting the feature of many real-world data-driven optimization problems. Experimental results show the benefits (and also cost) of our proposed DES meta-framework for large-scale BBO (see Section IV for details). In principle, this meta-framework can also be extended to other evolutionary algorithms with modifications.
@@ II Related Works
In this section, we review only parallel/distributed versions and large-scale variants of ES, since there have been some well-written reviews for ES (e.g., [35, 21, 50, 51, 52, 53, 54]) up to now.
@@ II-A Parallel/Distributed Evolution Strategies
Recent advances in parallel/distributed computing, particularly cloud computing [55, 56, 57], provide new advantages and challenges for evolutionary algorithms (EAs) [12, 58]. Although it comes as no surprise that parallelism is not a panacea for all cases [59, 60, 61] distributed evolutionary strategies (DES) are playing an increasingly important role in large-scale BBO in the post-Moore era [6, 62]. Refer to e.g., the recently proposed hardware lottery [63] for interesting discussions. Note that here we focus on model-level or application-level parallelism222Although some researchers viewed some distributed EAs as a new class of meta-heuristics, in this paper we adopt a conservative perspective, that is, distributed EAs are seen as a performance enhancement under distributed computing. (rather than instruction-level parallelism [7]).
According to [64], one of the first to parallelize ESs is [65], where an outdated vector computer was employed in 1983. In the early days of ES research, Schwefel [43] used the classical master-slave [66, 67, 68] model to conduct evolutionary (collective) learning of variable scalings on parallel architectures333The master-slave model (aka farmer/worker [69, 70]) is typically used for computationally-intensive fitness evaluations such as optimization of aircraft side rudder and racing car rear wing [71], etc. The well-established Amdahl’s law [72, 73] can be used for an often useful speedup estimation.. However, only a simulated (not realistic) parallel environment was used in his experiments, where the costs of data communication, task scheduling, and distributed fault-tolerance were totally ignored. It may over-estimate the convergence performance of DES. This issue existed mainly in early DES studies such as [74, 75, 76] given the fact that at that time commercial parallel/distributed computers were not widely available and Moore’s law worked well.
@@
Wilson et al. [85] proposed an asynchronous communication protocol to parallelize the powerful CMA-ES on cloud computing platforms. When the original CMA-ES was used as the basic computing unit for each CPU core, however, under its quadratic computational complexity the problem dimensions to be optimized are often much low (e.g., only 50 in their paper) by the modern standard for large-scale BBO. Similar issues are also found in existing libraries such as pCMALib [86], Playdoh [87], OpenFPM [88] and [89]. Glasmachers [90] updated strategy parameters asynchronously for Natural ES (NES) [91], a more principal version for ES. The runtime speedup ratio obtained was below 60% on the 8-d Rosenbrock function, which means the need of improvements. Reverse and Jaeger [92] designed a parallel island-based ES called piES to optimize a non-linear (62-d) systems biology model [93]. In piES, only individual step-sizes were self-adapted for each island (the CMA mechanism was ignored). The speedup formulation used in their paper considered only the runtime but not the solution quality, which may lead to over-optimistic conclusions in many cases.
Recently, Cuccu et al. [94] proposed a DES framework called DiBB based on the partially separable assumption [95, 96]. Although it obtained significant speedups when this assumption is satisfied, DiBB did not show obvious advantages than the serial CMA-ES on ill-conditioned non-separable landscapes [94]. Kucharavy et al. [97] designed a Byzantine-resilient [98] consensus strategy for DES. However, they did not explicitly consider the improvement of search performance and no performance data was reported in their paper. Rehbach et al. [99] used the classical (1+1)-ES as the local searcher for parallel model-based optimization on a small-sized (i.e., 16-core) parallel hardware. To our knowledge, they did not consider the more challenging distributed computing scenarios.
Another (perhaps less-known) research line for DES is Meta-ES (also referred to as Nested-ES, first proposed by Rechenberg [100, 101]) which organizes multiple independent (parallel) ESs hierarchically [39, 102]. Although there have been relatively rich theoretical works on different landscapes (e.g., parabolic ridge [103], sphere [104], noisy sphere [40], sharp ridge [105], ellipsoid [106], conic constraining [107]), to our knowledge all these theoretical models do not take overheads from distributed computing [49] into account. Furthermore, although these works provide valuable theoretical insights to understand Meta-ES, all the inner-ESs used in their models are relatively simple from a practical viewpoint.
@@ II-B Large-Scale Variants of CMA-ES
In this subsection, we review large-scale variants of CMA-ES through the lens of distributed computing. For a detailed introduction, refer to e.g., [113, 114, 115, 24] and references therein.
Because the standard CMA-ES has a quadratic time-space complexity, it is difficult to directly distribute it on cloud or clustering computing platforms for large-scale BBO. A key point to alleviate this issue is to reduce the computational complexity of the CMA mechanism, in order to fit better for the (distributed) memory hierarchy. Till now, different ways have been proposed to improve its computational efficiency: 1) exploiting the low-rank structure, e.g., [114, 115, 116, 117, 118]; 2) making the separable assumption, e.g., [119, 120]; 3) inspiring from L-BFGS, e.g. [19, 121, 20]; 4) seeking computationally more efficient implementations [122, 123, 124, 125, 126, 127, 128].
@@
One key challenge for DES lies in the trade-off between (computation) simplicity and (model) flexibility. On the one hand, we need to keep the structure of CMA, which can be simply parameterized as the number of evolution paths to reconstruct, as simple as possible, in order to fit better for memory hierarchy and reduce communication costs. On the another hand, we also expect to maintain the well-established invariance property as much as possible, in order to keep the flexible expressiveness/richness of model (resembling the second-order optimization method) [129]. To achieve such a trade-off, an efficient adaptive strategy (particularly at the meta-level) is highly desirable, which is the goal of our paper.
@@ III A Multilevel Meta-Framework for DES
In this paper we propose a multilevel-based meta-framework for distributed evolution strategies (DES) to minimize444Without loss of generality, the maximization problem can be easily transferred to the minimization problem by simply negating it. the large-scale BBO problem , where is the dimensionality. Very recently, a research team from the biologist Koonin has presented a mathematical model of evolution through the lens of multilevel learning in their two PNAS papers [37],[38]. Inspired by this evolution theory, our meta-framework for DES mainly involves hierarchical organization of distributed computing units (via Meta-ES), multilevel selection555For the modern theory regarding to evolutionary transitions in individuality [130],[131], multilevel selection is regarded as a crucial factor to understand life’s complexification [132],[133]., and collective learning of parameterized search/mutation distributions on structured populations [134]. While this new evolution theory is interpreted in a mathematical way, our meta-framework needs to be well-interpreted from an optimizer view of point, as shown below in detail.

### v4

@@ Abstract
In the post-Moore era, main performance gains of black-box optimizers are increasingly depending on parallelism, especially for large-scale optimization (LSO). Here we propose to parallelize the well-established covariance matrix adaptation evolution strategy (CMA-ES) and in particular its one latest LSO variant called limited-memory CMA-ES (LM-CMA). To achieve efficiency while approximating its powerful invariance property, we present a multilevel learning-based meta-framework for distributed LM-CMA. Owing to its hierarchically organized structure, Meta-ES is well-suited to implement our distributed meta-framework, wherein the outer-ES controls strategy parameters while all parallel inner-ESs run the serial LM-CMA with different settings. For the distribution mean update of the outer-ES, both the elitist and multi-recombination strategy are used in parallel to avoid stagnation and regression, respectively. To exploit spatiotemporal information, the global step-size adaptation combines Meta-ES with the parallel cumulative step-size adaptation. After each isolation time, our meta-framework employs both the structure and parameter learning strategy to combine aligned evolution paths for CMA reconstruction. Experiments on a set of large-scale benchmarking functions with memory-intensive evaluations, arguably reflecting many data-driven optimization problems, validate the benefits (e.g., effectiveness w.r.t. solution quality, and adaptability w.r.t. second-order learning) and costs of our meta-framework.
@@
Specifically, we consider the derandomized evolution strategy with covariance matrix adaptation (CMA-ES [16, 17, 18]) and in particular its one latest variant called limited-memory CMA (LM-CMA [19, 20]) for large-scale black-box optimization (BBO). As stated in the popular Nature review [12], “CMA-ES is widely regarded as (one of) the state of the art in numerical (black-box) optimization” with competitive performance on many benchmarking functions [21, 22, 23, 24, 25] and challenging applications (such as [26, 27, 28, 29, 30, 31], just to name a few). Although typically its absolute runtime is of minor relevance [17] for low-dimensional (e.g., 50) cases, it cannot be ignored in the distributed computing context for large-scale (e.g., 1000) optimization (LSO) because of its (at least) quadratic complexity w.r.t. each sampling. Since its limited-memory variant (LM-CMA) can fit better for memory hierarchy and distributed communication, our aim is to extend LM-CMA to the modern cloud/clustering computing environment for LSO, in order to show effectiveness [32] while approximating the attractive invariance property of CMA-ES [33] as much as possible.
@@
In their seminal paper [17], Hansen and Ostermeier pointed out four fundamental demands for ESs: adaptation, performance, invariance [34], and stationarity (i.e., unbiasedness under random/neutral selection). Clearly, these demands111As was previously stated in the classical ES review by Beyer and Schwefel [35], if some principle is deviated when we design the ES-based variant, the designed optimizer often needs to be widely tested. are also highly desirable for any distributed ES (DES), in order to obtain efficiency and generalization/transferability [36]. To meet these demands in the distributed computing environment, we adopt the multilevel learning perspective for evolution (recently published in PNAS [37, 38]) to model and design the efficient DES framework, wherein Meta-ES [39, 40] and LM-CMA could be naturally combined together to enjoy the best of both worlds.
With the rise of deep models and big data, currently there are increasing needs to optimize high-dimensional objective functions. Among them, a number of black-box scenarios from e.g., non-differentiable simulations and non-convex mathematical models urgently require black-box optimizers to obtain satisfactory performance in a reasonable runtime. Given the fact that almost all of serial black-box optimizers easily suffer from the curse-of-dimensionality issue, parallelism is a very natural way to scale up them to large-scale optimization, which is the focus of our paper. Three main contributions of our paper to DES for large-scale BBO are presented as below:
1) We first analyze the parallelism opportunities and challenges of two different ES families (i.e., CMA-ES and Meta-ES) under two common models (i.e., coordinator-worker and island). See Section II for details. We argue that multilevel learning for biological evolution (MLE) is a natural way to hierarchically combine LM-CMA with Meta-ES under distributed computing (Section III).
@@
2) Inspired by MLE [37], we propose a multilevel learning-based meta-framework for DES to exploit spatio-temporal information (if available) on-the-fly to accelerate convergence while maintaining meta-population diversity. Within it, four following key design choices for DES are made reasonably, in order to balance search efficiency (w.r.t. convergence rate, absolute runtime, and diversity) and extra computing cost brought by distributed enhancements (e.g., distributed scheduling, load balancing, fault-tolerance, and data exchanges over the network, etc.) [41].
@@ II Related Works
ES Version | Assumption | Complexity | Model Expressability | Challenge for LSO |
DiBB[49] | partial separability | high | weakness on non-separable landscapes | |
piES[50] | non-separability | low | adaptation of only individual step-sizes | |
asynchronous NES[51] | non-separability | very high | excessive requirements in CPU memory | |
asynchronous CMAES[52] | non-separability | high | difficulty in analyzing procedures | |
distributed ES[53] | island | low | adaptation of only the global step-size | |
[…]
VDCMA[55], MMES[56] | fixed low-rank | modest | no parallelism | |
R1ES[57], R1NES[58] | predominated direction | low | no parallelism | |
DDCMA[57] | separability & invariance | high | no parallelism | |
SEPCMAES[59], SNES[58] | separability | low | no parallelism | |
MAES[60], CCMAES[61] | invariance | high | no parallelism |
: is the dimensionality of problem to be optimized.
: is the total number of evolution paths used to reconstruct the covariance matrix (typically ).
: is the maximal dimensionality of all subspaces after decision space decomposition.
In this section, we only review parallel/distributed versions and large-scale variants of ES, since there have been some well-written reviews for ES (e.g., [35, 21, 62, 63, 64, 65]) up to now. We also analyze the opportunities and challenges of two different ES families (i.e., CMA-ES and Meta-ES) under two parallelism models (i.e., coordinator-worker and island).
@@ II-A Parallel/Distributed Evolution Strategies
Recent advances in parallel/distributed computing, particularly cloud computing [66, 67, 68], provide new advantages and challenges for evolutionary algorithms (EAs) [12, 69]. Although it comes as no surprise that parallelism is not a panacea for all cases [70, 71, 72], DES are playing an increasingly important role in large-scale BBO in the post-Moore era [6, 73]. Refer to e.g., the recently proposed hardware lottery [74] for insightful discussions. Note that here we focus on only model-level or application-level parallelism222Although some researchers viewed some distributed EAs as a new class of meta-heuristics, here we adopt a conservative perspective, that is, distributed EAs are seen as a performance enhancement under distributed computing. (rather than instruction-level parallelism [7]).
According to [75], one of the first to parallelize ESs was [76], where an outdated vector computer was employed in 1983. In the early days of ES research, Schwefel [77] used the classical coordinator-worker [78, 79, 80] model to conduct evolutionary (collective) learning of variable scalings on parallel architectures333The coordinator-worker model (aka farmer/worker [81, 82]) is typically used for computationally-intensive fitness evaluations such as optimization of aircraft side rudder and racing car rear wing [81]), etc. The well-established Amdahl’s law [83, 84] can be used as an often useful speedup estimation.. However, only a simulated (not realistic) parallel environment was used in his experiments, where the costs of data communication, task scheduling, and distributed fault-tolerance were totally ignored. It may over-estimate the convergence performance of DES. This issue existed mainly in early DES studies such as [85, 53, 86] given the fact that at that time commercial parallel/distributed computers were not widely available and Moore’s law still worked well.
@@
Wilson et al. [52] proposed an asynchronous communication protocol to parallelize the powerful CMA-ES on cloud computing platforms. When the original CMA-ES was used as the basic computing unit for each CPU core, however, under its quadratic computational complexity the problem dimensions to be optimized are often much low (e.g., only 50 in their paper) by the modern standard for large-scale BBO. Similar issues are also found in existing libraries such as pCMALib [95], Playdoh [96], OpenFPM [97] and [98]. Glasmachers [51] updated strategy parameters asynchronously for Natural ES (NES) [99], a more principal version for ES. The runtime speedup ratio obtained was below 60% on the 8-d Rosenbrock function, which indicates the need for improvements. Reverse and Jaeger [50] designed a parallel island-based ES called piES to optimize a non-linear (62-d) systems biology model [100]. In piES, only individual step-sizes were self-adapted for each island (the CMA mechanism was ignored). The speedup formulation used in their paper considered only the runtime but not the solution quality, which may lead to over-optimistic conclusions in many cases.
Recently, Cuccu et al. [49] proposed a DES framework called DiBB based on the partially separable assumption [101, 102]. Although it obtained significant speedups when this assumption was satisfied, DiBB did not show obvious advantages against the serial CMA-ES on ill-conditioned non-separable landscapes [49]. Kucharavy et al. [103] designed a Byzantine-resilient [104] consensus strategy for DES. However, they did not explicitly consider the improvement of search performance and no performance data was reported in their paper. Rehbach et al. [105] used the classical (1+1)-ES as the local searcher for parallel model-based optimization on a small-sized (i.e., 16-core) parallel hardware. To our knowledge, they did not consider the more challenging distributed computing scenarios.
Another (perhaps less-known) research line for DES is Meta-ES (also referred to as Nested-ES, first proposed by Rechenberg [106, 107]) which organizes multiple independent (parallel) ESs hierarchically [39, 108]. Although there have been relatively rich theoretical works on different landscapes (e.g., parabolic ridge [109], sphere [110], noisy sphere [40], sharp ridge [111], ellipsoid [112], conic constraining [113]), to our knowledge, all these theoretical models do not take overheads from distributed computing [48] into account. Furthermore, although these works provide valuable theoretical insights to understand Meta-ES, all the inner-ESs used in their models are relatively simple from a practical viewpoint.
@@ II-B Large-Scale Variants of CMA-ES
In this subsection, we review large-scale variants of CMA-ES through the lens of distributed computing. For an introduction, refer to e.g., [119, 54, 55, 24] and references therein.
Because the standard CMA-ES has a quadratic time-space complexity, it is difficult to directly distribute it on cloud or clustering computing platforms for large-scale BBO. A key point to alleviate this issue is to reduce the computational complexity of the CMA mechanism, in order to fit better the (distributed) memory hierarchy. Till now, different ways have been proposed to improve its computational efficiency: 1) exploiting the low-rank structure, e.g., [54, 55, 57, 120, 58]; 2) making the separable assumption, e.g., [121, 59]; 3) inspiring from L-BFGS, e.g. [19, 56, 20]; 4) seeking computationally more efficient implementations [60, 122, 61, 123, 124, 125, 126].
@@
One key challenge for DES lies in the trade-off between (computation) simplicity and (model) flexibility. On the one hand, we need to keep the structure of CMA, which can be simply parameterized as the number of evolution paths to reconstruct, as simple as possible, in order to fit better memory hierarchy and reduce communication costs. On the another hand, we also expect to maintain the well-established invariance property as much as possible, in order to keep the flexible expressiveness/richness of model (resembling the second-order optimization method) [127]. To achieve such a trade-off, an efficient adaptive strategy (particularly at the meta-level) is highly desirable, which is the goal of our paper.
Our paper will enhance the state of the art from three main aspects: 1) more powerful adaptation strategies under the distributed framework as compared with existing parallel/distributed ES versions (e.g., [49, 51, 52]); 2) efficient utilization of parallel/distributed computing resources when compared against serial large-scale ES versions (e.g., [55, 58, 60]); 3) an open-source implementation on a relatively large clustering computing platform, which is really rare in the current DES research, to our knowledge.
@@ III A Multilevel Meta-Framework for DES
In this paper we propose a multilevel-based meta-framework for distributed evolution strategies (DES) to minimize444Without loss of generality, the maximization problem can be easily transferred to the minimization problem by simply negating it. the large-scale BBO problem , where is the dimensionality. Very recently, a research team led by the biologist Koonin has presented a mathematical model of evolution through the lens of multilevel learning in their two PNAS papers [37],[38]. Inspired by this evolution theory, our meta-framework for DES mainly involves hierarchical organization of distributed computing units (via Meta-ES), multilevel selection555For the modern theory regarding to evolutionary transitions in individuality [128],[129], multilevel selection is regarded as a crucial factor to understand life’s complexification [130],[131]., and collective learning of parameterized search/mutation distributions on structured populations [132]. While this new evolution theory is interpreted in a mathematical way, our meta-framework, whose flowchart is shown at an relatively abstract level in Fig. 1, needs to be well-interpreted from an optimizer view of point, as shown below in detail.